OpenStudy (anonymous):

Write an equation of a line that passes through the point (8, 4) and is parallel to the line y = 4x + 2. y = 4x − 28 y = 4x + 28 y = 1 over 4x − 2 y = 1 over 4x + 2

OpenStudy (nirmalnema):

2 parallel lines have same slope.. so slope of given line is -a/b=-(4/-1)=4 now use one point formula y-y1=m(x-x1) y-4=4(x-8) y-4=4x-32 4x-y-28=0 y=4x-28

OpenStudy (anonymous):

Thank you for the explanation also, it cleared it up for me. I would feel bad if I ask for more, which I only have two. But thank you!

OpenStudy (nirmalnema):

no u can ask no problem in that

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

I don't want to take up your time.

OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at C (−1, 2), D (−2, −1), E (2, −2), and F (1, 1). Round your answer to the nearest hundredth when necessary. 10.92 units 12.68 units 13.43 units 18.27 units --- I have tried this multiple times and I can't seem to get it. Last time I got 7....(decimals), which is not a answer.

OpenStudy (nirmalnema):

do u need full explanation or simply i give the ans

OpenStudy (anonymous):

You can just give the answer.

OpenStudy (nirmalnema):

u can find the side lenghts by distance formula \[\sqrt{(x1-x2)^2+(y1-y2)^2}\]

OpenStudy (anonymous):

I will give that a try. I hope it works, do you mind a few more?

OpenStudy (nirmalnema):

yaa

OpenStudy (anonymous):

What relationship do the ratios of sin x° and cos y° share? The ratios are opposites (16 / 20 and -16 / 20) The ratios are both identical (16 / 20 and 16 / 20) The ratios are both negative (-16 / 20 and -16 / 20) The ratios are reciprocals (16 / 20 and 20 / 16)

OpenStudy (nirmalnema):

cos=base / hyp sin=per / hyp The ratios are both identical (16 / 20 and 16 / 20)

OpenStudy (anonymous):

Thank you, this is my last one. Its another word one and I have never done good with these questions.

OpenStudy (anonymous):

Line BC has an equation of a line y = 2x + 3, and line EF has an equation of a line y = −3x + 4. These two equations represent: parallel lines because the slopes of the lines are the same perpendicular lines because the product of the slopes is −1 the same line lines that are neither parallel nor perpendicular

OpenStudy (nirmalnema):

lines that are neither parallel nor perpendicular

OpenStudy (anonymous):

Thank you so much, you were a big help!

OpenStudy (nirmalnema):

your welcome